Engineering Tasks for UT 06/07 April 2019

  1. LBC:
    • ALERT! Guiding issue - rotational trail/drift as a function of dither position (or not) https://lbt.issuetrak.com/CSIssue_View.asp?IssueNbr=4450
      • Replicate observations (target, elevation, dither/no dither) of observation that showed this...most recent UT 2019-03-31 (see plot attached to IT 4450)
      • POSTPONED - Not sure if replicating this error tells us anything
  2. LUCI - Dithering test - seeing < 1.25" (DIMM)
    • SL-Bino
    • ESM-DX
      • The scripts are in /home/lbto/Service_Observing/E_dithertest
        • E_Feb17UT_dithertest.xml (OT program file)
        • Scripts_BinoSL_06to12hr/
        • Scripts_BinoSL+L2ESM_06h54d/
        • Scripts_BinoSL+L2ESM_14h10d/ <== field available during early April 2nd half E nights
  3. AGW Transforms LUCI
    • Note: different MOS masks in each LUCI (L2 exchange 4/5, L1 exchange 4/8) - there are a few INAF masks common to both exchanges
  4. DONE Software Group: https://wiki.lbto.org/Software/ENightWorkRequests
    • Tests are transparent to telescope operations - no special requests or requirements
    • errors occrred, reverted TCS and MCSPU
  5. Poor Weather Backups:
    • LUCI-Bino H-band Library (M-stars and test QSO target)
  6. ESM on-sky tests - (support from D. Miller and/or G. Taylor needed)
    • SOUL rollover if SOUL team can't do in first half
      • AO Ref star with R=8, 12, 16.5, 17.0, 17.5 - test of how faint an AO Ref star can be used
      • Stability testing - 30 min in a field, dithering - Target TBD
    • DX-ESM / SX-SL Tests
      • M92 - add additonal data points - and include H and J - test R=9, 15 AO Ref stars
      • Polar Ring Galaxy - AO Ref Star = 13.4 = Nucleus of Galaxy - test of using nucleus as AO Ref star
      • Hickson Compact Group Galaxy test - measure the change as a distance of AO Ref star for resolved objects as a function of distance
  7. Non-Sidereal guiding tests
  8. DONE OAC test - test preset and see if guidestar is recognized correctly - tested by reverting to old TCS and no error
  9. Test WFS for NS objects - reduce the WFS update to 4sec then 1sec to see if it keeps up with guidetar duing NS motion

Nightlog UT Apr 06, 2019 - 1st half SOUL Testing / 2nd half LBTO Technical/Eng

OSA: D.G. Huerta

ISp: J. Morris

ISA: B. Rothberg (Tucson)

AO: D. Miller (Tucson)

Featured Players: S. Esposito, E. Pinna, J. Christou, D. Miller, J. Hill (Tucson)

UT 01:48 - clouds visiile but clearing - currently experiencing TCS problems

UT 02:30 - Still working on TCS and various sub-systems - clouds look better again

UT 02:38 - TCS and sub-systems back up - init all on LUCIs

UT 03:14 - MCSPU issues - still not open

UT 03:49 - test preset (closed dome) worked OK

UT 03:59 - enclosure open - collimating on SX

UT 04:08 - took some time to collimate - moving to first AO star

UT 04:22 - had to stop and restart AOS sub-system - had some issues with PA of guidestar initially - finally got a good preset, but suddenly lost the guidestar and lost elevation information

-MCSPU needs to be restarted - somehow rotation stopped - which lost elevation info and stopped rotation which caused gs to move out of field

UT 04:30 - restarting AdSec - lost info on GUIs

UT 04:34 - sending new preset

-seems to be working - some adjustments needed to filter and sub-frame readouts


> OAC issue? when sending a manual preset, had to put PA=180 to get the GS into the field.

UT 05:32 - MCSPU died

UT 05:34 - reverting back to old MCSPU and old TCS

-this will likely take about 30minutes

UT 05:51 - reconnecting everything


> OAC test - sent a PA=180 and got an error, this suggests the new TCS upgarde had an error - so all is well now with old version (and much debugging to be done)

UT 06:15 - testing continuing - noting that the Humidity has been creeping up slowly since we started - at 83% now

UT 07:28 - humidity at 90.5%

UT 08:07 - handing over to ENG LBTO

UT 08:08 - going to start with Olga's dither test L2 ESM

UT 08:14 - reauthorizing for LUCI-2 only

UT 08 :22 - hit 95% humidity - closing

UT 10:32 - stil closed Humidity currently 87.7% - has dipped below 90% a few times in the last hour - clouds overhead

UT 10:47 - Humidity back up over 90%

UT 11:12 - back down to 87% humidity

UT 11:15 - Humidity is 80.5%

UT 11:20 - Humidity is 82.4%

UT 11:27 - hit 96.1% humidity

UT 11:56 - DGH Calling night

-closing out GUIs, LUCIs in safe state.


-- BarryRothberg - 02 Apr 2019
Topic revision: r13 - 06 Apr 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback